Fermi National Laboratory
 

Operation Summary for September 21, 2009 - September 23, 2009

 

Monday Sept. 21

The day shift began with Operations monitoring the Tevatron (TeV) store 7196 (82.49E30) and the Recycler stash (174.5E10), with the Antiproton Source stacking (23.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI experiment taking beam, with MiniBooNE off due to heating, ventilation, and air conditioning (HVAC) work, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations terminated store 7196 at 9:01 AM. TeV personnel began studies.

Operations turned off the Pelletron at 9:45 AM. This was to investigate controls and regulation problems, and a critical device controller PLC problem, which was preventing transfers of antiprotons. The Pelletron came back on line at 3:53 PM.

Operations put the TeV into shot setup and established store 7201 at 7:15 PM with an average initial luminosity (AIL) of 204.44E30.

Tuesday Sept. 22

The midnight shift began with Operations monitoring store 7201 (106.36E30) and the stash (120.3E10), with Pbar stacking (17.9E10), with NuMI taking beam, and with MiniBooNE still off for HVAC work.

Operators completed a search and secure of the Switchyard and Meson enclosures by 4:56 AM for safety system testing on the Tuesday day shift.

Vacuum bursts in the MI-30 area persisted throughout the day and evening shifts.

Operations terminated store 7201 at 12:03 PM.

Operations put the TeV into shot setup and established store 7202 at 2:23 PM with an AIL of 200E30.

Operations established beam to MiniBooNE at 2:16 PM; it took beam throughout the evening shift while work on the HVAC continued.

A MI kicker (KPS5S) tripped off at 2:52 PM due to a coolant leak and a bad high voltage connector. MI personnel completed these repairs and resumed sending beam to Pbar by the end of the midnight shift.

Wednesday Sept. 23

The midnight shift began with Operations monitoring store 7202 (73.13E30) and the stash (9.1E10), with Pbar stacking (47.5E10), and with NuMI and MiniBooNE taking beam.

Vacuum bursts in the MI-30 area continued throughout the midnight shift.

HVAC work in MiniBooNE stopped at 1:25 AM. It will continue on the Wednesday day shift. Operations turned off the MiniBooNE beam.

There were some short downtimes throughout the midnight shift due to Booster RF stations (BRF3, 6, and 15) and to a Linac RF station (LRF3).

The Plans for Wednesday

The plans for today are to establish a new store by mid morning. After the store, MI will access and investigate their vacuum and extraction septa problems. The Neutron Therapy Facility will treat patients until noon, and then Preaccelerator personnel will access the H- Source dome.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/23/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory